Shift summary, Evening, May 30th.

Robert Tschirhart (tsch@fnal.gov)
Mon, 31 May 1999 00:25:43 -0500

Shift Summary for {Evening} shift: May 30, 1999
On Shift: Doug Jensen - 1st half
Bob T - 2nd half

A variety of other committed collaborators were working throughout this
shift.
Theo and Ashkan continued work for about an hour - Slow DA
Peter Shanahan - worked on CsI stuff

Hours of beam delivered at all intensities: 0.0
Hours of beam delivered at nominal intensity: 0.0
Hours of 99 data (special runs not included): 0
Hours of special runs (laser scans, muon runs...): 0

Accounting of lost time:
No beam this shift due to TeV studies.

Activities:

Show up for shift. Status: No beam, running short shifts. Theo and
Ashkan finsihing up some SDAQ work, Peter Shanahan working on CsI

1645 Check Z's with laser ranger.
BP oil fine. NM3 remains interlocked.
1710 Do Gas Check list

Peter finished work on CsI - Change L1 delay from 108 to 105 ns.
This fixes the capid problems. No more CsI errors. He leaves about
1740 pm.

1820 Trigger error message from Elliot.
Reset/kfriend and it goes again. Thanx to Valerie for the help.


1840 Valerie leaves, having tweeked up the VV' gains.

1850 Checked tiltmeters again. The Nivel Tiltmeter in the CsI house
is not reading back. I have checked the cables outside the
blockhouse. Next time somebody is in the blockhouse, could they
please check that the cable is plugged in and that things look
OK with the tiltmeter? Let me know when it is done ( would
like to join the entry if that works out ). djensen@fnal.gov.


1900 New 1/2 shift takes over. Inherited Cosmics run 13237, which
continues to run smoothly.

2100 Miscellaneous inspection of the hall turned up the following
problems:

Bad or dead fans in need of replacement:

Upstairs: Rack RR180. FERA.
Rack RR167. Trigger.

Downstairs: Rack RR101. Muon system.
Rack RR042. SA2 Shaper fan, flakey.

Found that CsI rib temperatures 13-T and 15-T were not
reading correctly. Connector reseated, now fine...except
that this act has caused two CsI channels to jump in gain
by 2-3%.

Noticed that SDAQ is reporting that one of the V-prime
counters has a HV out of range. This is probably due to the
HV adjustments that were made. Reference should be updated
in SDAQ accordingly.

23:50 On arrival of new shift crew, DAQ hangs with permit problems
in stream-5. More news as it breaks.

=========================================================================
Runs ended this shift:
Run 13237 running - no tape

Tapes brought to FCC this shift: 0